### v1

# Resolvent estimates in Schatten spaces for Laplace-Beltrami operators on compact manifolds

Abstract.
We prove resolvent estimates in Schatten spaces for Laplace-Beltrami operators on compact manifolds at the critical exponent. Our proof only uses known bounds for the Hadamard parametrix.
1. Introduction and main result
Let $(M,g)$ be a compact boundaryless Riemannian manifold of dimension $n\geq 2$, with smooth metric $g$, and consider the negative Laplace–Beltrami operator $\Delta_{g}$ on $M$. The aim of this note is to prove a Schatten space version of the following resolvent estimate,
(1) $\|(-\Delta_{g}-z)^{-1}\|_{L^{q^{\prime}}(M)\to L^{q}(M)}\leq C_{\delta}|z|^{\sigma(q)-\frac{1}{2}},$
where $\mathrm{Im}\,\sqrt{z}\geq\delta$ and
(2) $\sigma(q)=\max\bigl{(}\,n(\tfrac{1}{2}-\tfrac{1}{q})-\tfrac{1}{2},\,\tfrac{n-1}{2}(\tfrac{1}{2}-\tfrac{1}{q})\,\bigr{)}.$
Here, $\delta>0$ is an arbitrary but fixed constant. For $n\geq 3$, $q=2n/(n-2)$, inequality (1) was proved by Dos Santos Ferreira–Kenig–Salo [6]. For the torus this result is due to Shen [15]. Bourgain–Shao–Sogge–Yao [2] obtained improved bounds on negatively curved manifolds and on the torus. Frank–Schimmer [7], and, independently, Burq–Dos Santos Ferreira–Krupchyk [3], proved (1) for $n\geq 2$ at the the critical exponent $q_{n}:=2(n+1)/(n-1)$. By elliptic estimates, Sobolev embedding and interpolation with the trivial $L^{2}\to L^{2}$ bound, the estimate at the critical exponent implies that (1) holds for all other exponents $2\leq q\leq 2n/(n-2)$. We will therefore only consider the critical exponent in the following.
By duality, the inequality (1) at the critical exponent is equivalent to
(3) $\|W_{1}(-\Delta_{g}-z)^{-1}W_{2}\|_{L^{2}(M)\to L^{2}(M)}\leq C_{\delta}|z|^{\frac{1}{q_{n}}-\frac{1}{2}}\|W_{1}\|_{L^{n+1}(M)}\|W_{2}\|_{L^{n+1}(M)},$
where we used that $\sigma(q_{n})=1/q_{n}$. Our main result is the following upgrade of (3) to a stronger bound that replaces the $L^{2}(M)\to L^{2}(M)$ operator norm by a Schatten norm.
Theorem 1.1.
Let $(M,g)$ be a compact boundaryless Riemannian manifold of dimension $n\geq 2$, with smooth metric $g$. Then for every $\delta>0$ there exists a constant $C_{\delta}$ such that
(4) $\|W_{1}(-\Delta_{g}-z)^{-1}W_{2}\|_{\mathfrak{S}^{n+1}(L^{2}(M))}\leq C_{\delta}|z|^{\frac{1}{q_{n}}-\frac{1}{2}}\|W_{1}\|_{L^{n+1}(M)}\|W_{2}\|_{L^{n+1}(M)}.$
Resolvent estimates in Schatten spaces for the Laplace–Beltrami operator on nontrapping asymptotically conic manifolds were established by Guillarmou–Hassell–Krupchyk [10]. The proof for compact manifolds presented here is much shorter.
Our resolvent estimate is closely related to the following spectral cluster bound of Frank–Sabin [8] (this is the estimate in the form (16) there),
(5) $\|W\Pi_{\lambda}\overline{W}\|_{\mathfrak{S}^{n+1}(L^{2}(M))}\lesssim(1+\lambda)^{\frac{2}{q_{n}}}\|W\|^{2}_{L^{n+1}(M)}.$
Here $\Pi_{\lambda}:=\mathbf{1}_{[\lambda,\lambda+1]}(\sqrt{-\Delta_{g}})$ is the spectral projector onto frequencies in the unit length window $[\lambda,\lambda+1]$ for $\lambda>0$. We recall that a compact operator $K$ belongs to the the Schatten space $\mathfrak{S}^{\alpha}$ if
$\|K\|_{\mathfrak{S}^{\alpha}}:=\big{(}\sum_{j=1}^{\infty}s_{j}(K)^{\alpha}\big{)}^{\frac{1}{\alpha}}<\infty,$
where $s_{j}(K)$ are the singular values of $K$ (i.e. the eigenvalues of $(KK^{*})^{\frac{1}{2}}$). The dual estimate to (5) takes the form
(6) $\big{\|}\sum_{j\in J}\nu_{j}|f_{j}|^{2}\big{\|}_{L^{q_{n}/2}(M)}\lesssim(1+\lambda)^{\frac{2}{q_{n}}}\big{(}\sum_{j\in J}|\nu_{j}|^{\frac{n+1}{n}}\big{)}^{\frac{n}{n+1}},$
whenever $(f_{j})_{j\in J}\subset\mathop{\rm Ran}\Pi_{\lambda}$ is an orthonormal family of functions. In the case of a single function, (6) recovers Sogge’s spectral cluster bounds [16] at the critical exponent. One can interpolate (5) with the sharp counting function remainder estimate of Avakumović, Levitan and Hörmander [1, 14, 11],
$\operatorname{Tr}W\Pi_{\lambda}\overline{W}=\int_{M}|W(x)|^{2}\Pi_{\lambda}(x,x){\rm d}V_{g}\lesssim(1+\lambda)^{n-1}\|W\|_{L^{2}(M)}^{2},$
and with the trivial operator norm bound
$\|W\Pi_{\lambda}\overline{W}\|_{L^{2}(M)\to L^{2}(M)}\leq\|W\|^{2}_{L^{\infty}(M)}$
to obtain the full range of spectral cluster estimates
(7) $\big{\|}\sum_{j\in J}\nu_{j}|f_{j}|^{2}\big{\|}_{L^{q/2}(M)}\lesssim(1+\lambda)^{2\sigma(q)}\big{(}\sum_{j\in J}|\nu_{j}|^{\alpha(q)}\big{)}^{1/\alpha(q)},$
where
$\alpha(q):=\max\bigl{(}\,\tfrac{q(n-1)}{2n},\tfrac{2q}{q+2}\,\bigr{)}.$
Since $\mathrm{Im}\,(-\Delta_{g}-(\lambda+\mathrm{i})^{2})^{-1}\gtrsim\lambda^{-1}\Pi_{\lambda}$, the resolvent bound (4) implies (5) and thus (6), (7). On the other hand, it is not difficult to see that (5) implies (4) with a logarithmic loss, i.e. an additional factor of $\log(2+\lambda)$ on the right hand side (the proof is similar to that of [5, Prop. 3.3]). In [5], it was shown that the $L^{q^{\prime}}(M)\to L^{q}(M)$ resolvent estimates are actually a direct consequence of Sogge’s spectral cluster bounds. The idea of the proof is to apply the Christ–Kiselev lemma [4] to a microlocalized version of the spectral cluster bound. This strategy does not seem to work for Schatten norm bounds, so we give a direct proof of (4). We follow the approach of Dos Santos Ferreira–Kenig–Salo [6], which uses the Hadamard parametrix for the resolvent (see also Hörmander [12, 17.4] and Sogge [16]).
2. Hadamard parametrix
Following [6], let $T(z)$ be the operator with kernel
$T(x,y;z)=\chi(x,y)F(x,y;z),$
where $\chi$ is a localization to the diagonal $x=y$,
$F(x,y;z)=\sum_{\nu=0}^{N}a_{\nu}(x,y)F_{\nu}({\rm d}_{g}(x,y);z),$
with $N>(n-1)/2$, and
$F_{\nu}(|x|;z)=\nu!(2\pi)^{-n}\int_{{\mathbb{R}}^{n}}\frac{{\rm e}^{\mathrm{i}x\cdot\xi}}{(|\xi|^{2}-z)^{1+\nu}}{\rm d}\xi.$
The functions $a_{\nu}$ can be recursively chosen such that
(8) $(-\Delta_{g}-z)T(z)u=\chi(x,x)u+S(z)u,$
where the remainder $S(z)=S_{1}(z)+S_{2}(z)$ satisfies
(9) $S_{1}(x,y;z)=|z|^{\frac{n-1}{4}}{\rm e}^{-\sqrt{z}{\rm d}_{g}(x,y)}b(x,y;z),\quad S_{2}(x,y;z)=\mathcal{O}_{\delta}(|z|^{-1/2}),$
where $b$ is a smooth function, see (3.9) and the proof of Lemma 4.2 in [6].
As a technical device, we also define, for $0\leq\nu\leq N$ and $1+\nu+\mathrm{Re}\,s\in[0,(n+1)/2]$,
$F_{\nu+s}(|x|;z)=\Gamma(1+\nu+s)(2\pi)^{-n}\int_{{\mathbb{R}}^{n}}\frac{{\rm e}^{\mathrm{i}x\cdot\xi}}{(|\xi|^{2}-z)^{1+\nu+s}}{\rm d}\xi.$
These kernels define an analytic family of operators $s\mapsto F_{\nu+s}(z)$ which coincide with $F_{\nu}(z)$ for $s=0$.
Lemma 2.1.
For $|z|=1$, $z\neq 1$, we have
$\|F_{\frac{n-1}{2}+\mathrm{i}t}(z)\|_{L^{1}({\mathbb{R}}^{n})\to L^{\infty}({\mathbb{R}}^{n})}+\|F_{-1+\mathrm{i}t}(z)\|_{L^{2}({\mathbb{R}}^{n})\to L^{2}({\mathbb{R}}^{n})}\leq C{\rm e}^{c|t|^{2}}.$
Proof.
The bound for the first term can be found in [9, (50)] and is essentially contained in [13]. The bound for the second term is trivial. ∎
3. Proof of Theorem 1.1
Proof.
In the following, to shorten some of the estimates, we replace $W_{1}$ and $W_{2}$ with with $W$ and $\overline{W}$, respectively. Strictly speaking, this is weaker than the bound in Theorem 1.1 since the resolvent is not positive, but the proof for the general case is identical. We will also suppress the dependence of constants on $\delta$. By scaling, Lemma 2.1 and [9, Prop. 1] imply that (4) holds for $T(z)$. By (9), the second term $S_{2}(z)$ satisfies
$\|WS_{2}(z)\overline{W}\|_{\mathfrak{S}^{2}}\lesssim|z|^{-\frac{1}{2}}\|W\|^{2}_{L^{2}},$
which yields a bound that is better than (4) and hence can be ignored. By [8, Corollary 3] the first term satisfies
(10) $\|WS_{1}(z)\|_{\mathfrak{S}^{2(n+1)}}\lesssim|z|^{\frac{1}{2q_{n}}}\|W\|_{L^{n+1}}.$
After rescaling, Lemma 2.1 together with [9, Proposition 1] yields the parametrix bound
(11) $\|WT(z)\overline{W}\|_{\mathfrak{S}^{n+1}}\lesssim|z|^{\frac{1}{q_{n}}-\frac{1}{2}}\|W\|^{2}_{L^{n+1}},$
After summing (8) over a partition of unity we obtain (by slight abuse of notation)
(12) $(-\Delta_{g}-z)T(z)=\mathbf{1}+S(z).$
Note that, even though $T(z)$ and $S(z)$ are not symmetric, their adjoints satisfy the same bounds (10), (11). Moreover, these estimates remain true when replacing $z$ by $\overline{z}$. Using these observations and assuming, without loss of generality, that $\|W\|^{2}_{L^{n+1}(M)}=1$, we infer that
(13) $\|WT(z)\overline{W}\|_{\mathfrak{S}^{n+1}}+|z|^{-\frac{1}{2}}\|WS(z)^{*}\|_{\mathfrak{S}^{2(n+1)}}^{2}\lesssim|z|^{\frac{1}{q_{n}}-\frac{1}{2}},$
and we have the same bound with $\overline{z}$ in place of $z$. Applying the resolvent operator $(-\Delta_{g}-z)^{-1}$ to (12) yields
(14) $\|W(-\Delta_{g}-z)^{-1}\overline{W}\|_{\mathfrak{S}^{n+1}}\leq\|WT(z)\overline{W}\|_{\mathfrak{S}^{n+1}}+\|W(-\Delta_{g}-z)^{-1}S(z)\overline{W}\|_{\mathfrak{S}^{n+1}}.$
We first use the noncommutative Hölder inequality and (13) to bound
$\|W(-\Delta_{g}-z)^{-1}S(z)\overline{W}\|_{\mathfrak{S}^{n+1}}$ $\leq\|W(-\Delta_{g}-z)^{-1}\|_{\mathfrak{S}^{2(n+1)}}\|S(z)\overline{W}\|_{\mathfrak{S}^{2(n+1)}}$
$\lesssim|z|^{\frac{1}{2q_{n}}}\|W(-\Delta_{g}-z)^{-1}\|_{\mathfrak{S}^{2(n+1)}}.$
Then we use the resolvent identity
$(-\Delta_{g}-z)^{-1}(-\Delta_{g}-\overline{z})^{-1}=\frac{1}{2\mathrm{Im}\,z}((-\Delta_{g}-z)^{-1}-(-\Delta_{g}-\overline{z})^{-1})$
to obtain
$\|W(-\Delta_{g}-z)^{-1}\|_{\mathfrak{S}^{2(n+1)}}$ $=\|W(-\Delta_{g}-z)^{-1}(-\Delta_{g}-\overline{z})^{-1}\overline{W}\|_{\mathfrak{S}^{n+1}}^{1/2}$
$\lesssim|z|^{-\frac{1}{2}}(\|W(-\Delta_{g}-z)^{-1}\overline{W}\|_{\mathfrak{S}^{n+1}}+\|W(-\Delta_{g}-\overline{z})^{-1}\overline{W}\|_{\mathfrak{S}^{n+1}})^{1/2}.$
Combining the last two estimates and using that $ab\lesssim\epsilon a^{2}+\epsilon^{-1}b^{2}$ for arbitrary $\epsilon>0$, we get
$\|W(-\Delta_{g}-z)^{-1}S(z)\overline{W}\|_{\mathfrak{S}^{n+1}}$ $\lesssim\epsilon|z|^{-\frac{1}{2}}(\|W(-\Delta_{g}-z)^{-1}\overline{W}\|_{\mathfrak{S}^{n+1}}+\|W(-\Delta_{g}-\overline{z})^{-1}\overline{W}\|_{\mathfrak{S}^{n+1}})$
$+\epsilon^{-1}|z|^{\frac{1}{q_{n}}}$
We choose $\epsilon=c|z|^{\frac{1}{2}}$ for some sufficiently small constant $c$. Then, using (13), (14) and its analogue for $\overline{z}$, we can absorb the $\epsilon|z|^{-\frac{1}{2}}$ term to get
$\|W(-\Delta_{g}-z)^{-1}\overline{W}\|_{\mathfrak{S}^{n+1}}+\|W(-\Delta_{g}-\overline{z})^{-1}\overline{W}\|_{\mathfrak{S}^{n+1}}\lesssim|z|^{\frac{1}{q_{n}}-\frac{1}{2}}.$
This completes the proof of Theorem 1.1. ∎
Acknowledgements: The author thanks Rupert Frank, Mikko Salo, Jeff Galkowksi and Xiaoyan Su for useful discussions and correspondence. Support by the Engineering & Physical Sciences Research Council [grant number EP/X011488/1] is acknowledged.