### v2

# Resolvent estimates in Schatten spaces for Laplace-Beltrami operators on compact manifolds

Abstract.
We prove resolvent estimates in Schatten spaces for Laplace-Beltrami operators on compact manifolds at the critical exponent. Our proof only uses known bounds for the Hadamard parametrix.
1. Introduction and main result
Let $(M,g)$ be a compact boundaryless Riemannian manifold of dimension $n\geq 2$, with smooth metric $g$, and consider the negative Laplace–Beltrami operator $\Delta_{g}$ on $M$. The aim of this note is to prove a Schatten space version of the following resolvent estimate,
(1) $\|(-\Delta_{g}-z)^{-1}\|_{L^{q^{\prime}}(M)\to L^{q}(M)}\leq C_{\delta}|z|^{\sigma(q)-\frac{1}{2}},$
where $\mathrm{Im}\,\sqrt{z}\geq\delta$ and
(2) $\sigma(q)=\max\bigl(\,n(\tfrac{1}{2}-\tfrac{1}{q})-\tfrac{1}{2},\,\tfrac{n-1}{2}(\tfrac{1}{2}-\tfrac{1}{q})\,\bigr).$
Here, $\delta>0$ is an arbitrary but fixed constant. For $n\geq 3$, q=2n/(n2)q=2n/(n-2), inequality (1) was proved by Dos Santos Ferreira–Kenig–Salo [6]. For the torus this result is due to Shen [15]. Bourgain–Shao–Sogge–Yao [2] obtained improved bounds on negatively curved manifolds and on the torus. Frank–Schimmer [7], and, independently, Burq–Dos Santos Ferreira–Krupchyk [3], proved (1) for $n\geq 2$ at the critical exponent $q_{n}:=2(n+1)/(n-1)$. By elliptic estimates, Sobolev embedding and interpolation with the trivial $L^{2}\to L^{2}$ bound, the estimate at the critical exponent implies that (1) holds for all other exponents $2\leq q\leq 2n/(n-2)$. We will therefore only consider the critical exponent in the following.
By duality, the inequality (1) at the critical exponent is equivalent to
(3) $\|W_{1}(-\Delta_{g}-z)^{-1}W_{2}\|_{L^{2}(M)\to L^{2}(M)}\leq C_{\delta}|z|^{\frac{1}{q_{n}}-\frac{1}{2}}\|W_{1}\|_{L^{n+1}(M)}\|W_{2}\|_{L^{n+1}(M)},$
where we used that $\sigma(q_{n})=1/q_{n}$. Our main result is the following upgrade of (3) to a stronger bound that replaces the $L^{2}(M)\to L^{2}(M)$ operator norm by a Schatten norm.
Theorem 1.1.
Let $(M,g)$ be a compact boundaryless Riemannian manifold of dimension $n\geq 2$, with smooth metric $g$. Then for every $\delta>0$ there exists a constant $C_{\delta}$ such that
(4) $\|W_{1}(-\Delta_{g}-z)^{-1}W_{2}\|_{\mathfrak{S}^{n+1}(L^{2}(M))}\leq C_{\delta}|z|^{\frac{1}{q_{n}}-\frac{1}{2}}\|W_{1}\|_{L^{n+1}(M)}\|W_{2}\|_{L^{n+1}(M)}.$
Resolvent estimates in Schatten spaces for the Laplace–Beltrami operator on nontrapping asymptotically conic manifolds were established by Guillarmou–Hassell–Krupchyk [10]. The proof for compact manifolds presented here is much shorter.
Our resolvent estimate is closely related to the following spectral cluster bound of Frank–Sabin [8] (this is the estimate in the form (16) there),
(5) $\|W\Pi_{\lambda}\overline{W}\|_{\mathfrak{S}^{n+1}(L^{2}(M))}\lesssim(1+\lambda)^{\frac{2}{q_{n}}}\|W\|^{2}_{L^{n+1}(M)}.$
Here $\Pi_{\lambda}:=\mathbf{1}_{[\lambda,\lambda+1]}(\sqrt{-\Delta_{g}})$ is the spectral projector onto frequencies in the unit length window $\lambda,\lambda+1]$ for $\lambda>0$ and $A\lesssim B$ means $A\leq CB$ for some unspecified constant $C$. We recall that a compact operator $K$ belongs to the Schatten space $\mathfrak{S}^{\alpha}$ if
$\|K\|_{\mathfrak{S}^{\alpha}}:=\big(\sum_{j=1}^{\infty}s_{j}(K)^{\alpha}\big)^{\frac{1}{\alpha}}<\infty,$
where $s_{j}(K)$ are the singular values of $K$ (i.e. the eigenvalues of $(KK^{*})^{\frac{1}{2}}$). We refer to, e.g., [16] for background on Schatten spaces. The dual estimate to (5) takes the form
(6) $\big\|\sum_{j\in J}\nu_{j}|f_{j}|^{2}\big\|_{L^{q_{n}/2}(M)}\lesssim(1+\lambda)^{\frac{2}{q_{n}}}\big(\sum_{j\in J}|\nu_{j}|^{\frac{n+1}{n}}\big)^{\frac{n}{n+1}},$
whenever $(f_{j})_{j\in J}\subset\mathop{\rm Ran}\Pi_{\lambda}$ is an orthonormal family of functions. In the case of a single function, (6) recovers Sogge’s spectral cluster bounds [17] at the critical exponent. One can interpolate (5) with the sharp counting function remainder estimate of Avakumović, Levitan and Hörmander [1, 14, 11],
$\operatorname{Tr}W\Pi_{\lambda}\overline{W}=\int_{M}|W(x)|^{2}\Pi_{\lambda}(x,x){\rm d}V_{g}\lesssim(1+\lambda)^{n-1}\|W\|_{L^{2}(M)}^{2},$
and with the trivial operator norm bound
$\|W\Pi_{\lambda}\overline{W}\|_{L^{2}(M)\to L^{2}(M)}\leq\|W\|^{2}_{L^{\infty}(M)}$
to obtain the full range of spectral cluster estimates
(7) $\big\|\sum_{j\in J}\nu_{j}|f_{j}|^{2}\big\|_{L^{q/2}(M)}\lesssim(1+\lambda)^{2\sigma(q)}\big(\sum_{j\in J}|\nu_{j}|^{\alpha(q)}\big)^{1/\alpha(q)},$
where
$\alpha(q):=\max\bigl(\,\tfrac{q(n-1)}{2n},\tfrac{2q}{q+2}\,\bigr).$
Since $\mathrm{Im}\,(-\Delta_{g}-(\lambda+\mathrm{i})^{2})^{-1}\gtrsim\lambda^{-1}\Pi_{\lambda}$, the resolvent bound (4) implies (5) and thus (6), (7). On the other hand, it is not difficult to see that (5) implies (4) with a logarithmic loss, i.e. an additional factor of $\log(2+\lambda)$ on the right hand side (the proof is similar to that of [5, Prop. 3.3]). In [5], it was shown that the $L^{q^{\prime}}(M)\to L^{q}(M)$ resolvent estimates are actually a direct consequence of Sogge’s spectral cluster bounds. The idea of the proof is to apply the Christ–Kiselev lemma [4] to a microlocalized version of the spectral cluster bound. This strategy does not seem to work for Schatten norm bounds, so we give a direct proof of (4). We follow the approach of Dos Santos Ferreira–Kenig–Salo [6], which uses the Hadamard parametrix for the resolvent (see also Hörmander [12, 17.4] and Sogge [17]).
2. Hadamard parametrix
Following [6], let $T(z)$ be the operator with kernel
$T(x,y;z)=\chi(x,y)F(x,y;z),$
where $\chi$ is a localization to the diagonal $x=y$,
$F(x,y;z)=\sum_{\nu=0}^{N}a_{\nu}(x,y)F_{\nu}({\rm d}_{g}(x,y);z),$
with N>(n1)/2N>(n-1)/2, and
$F_{\nu}(|x|;z)=\nu!(2\pi)^{-n}\int_{{\mathbb{R}}^{n}}\frac{{\rm e}^{\mathrm{i}x\cdot\xi}}{(|\xi|^{2}-z)^{1+\nu}}{\rm d}\xi.$
The functions $a_{\nu}$ can be recursively chosen such that
(8) $(-\Delta_{g}-z)T(z)u=\chi(x,x)u+S(z)u,$
where the remainder $S(z)=S_{1}(z)+S_{2}(z)$ satisfies
(9) $S_{1}(x,y;z)=|z|^{\frac{n-1}{4}}{\rm e}^{-\sqrt{z}{\rm d}_{g}(x,y)}b(x,y;z),\quad S_{2}(x,y;z)=\mathcal{O}_{\delta}(|z|^{-1/2}),$
where $b$ is a smooth function, see (3.9) and the proof of Lemma 4.2 in [6]. As a technical device, we also define, for $0\leq\nu\leq N$ and $1+\nu+\mathrm{Re}\,s\in[0,(n+1)/2]$,
$F_{\nu+s}(|x|;z)=\Gamma(1+\nu+s)(2\pi)^{-n}\int_{{\mathbb{R}}^{n}}\frac{{\rm e}^{\mathrm{i}x\cdot\xi}}{(|\xi|^{2}-z)^{1+\nu+s}}{\rm d}\xi.$
These kernels define an analytic family of operators $s\mapsto F_{\nu+s}(z)$ which coincide with $F_{\nu}(z)$ for $s=0$.
Lemma 2.1.
For $|z|=1$, $z\neq 1$, we have
$\|F_{\frac{n-1}{2}+\mathrm{i}t}(z)\|_{L^{1}({\mathbb{R}}^{n})\to L^{\infty}({\mathbb{R}}^{n})}+\|F_{-1+\mathrm{i}t}(z)\|_{L^{2}({\mathbb{R}}^{n})\to L^{2}({\mathbb{R}}^{n})}\leq C{\rm e}^{c|t|^{2}}.$
Proof.
The bound for the first term can be found in [9, (50)] and is essentially contained in [13]. The bound for the second term is trivial. ∎
3. Proof of Theorem 1.1
Proof.
We first consider the case $W_{2}=\overline{W_{1}}$ and then prove the general case at the end. In the following, we suppress the dependence of constants on $\delta$. After rescaling, Lemma 2.1 together with [9, Proposition 1] yields the parametrix bound
(10) $\|W_{1}T(z)W_{2}\|_{\mathfrak{S}^{n+1}}\lesssim|z|^{\frac{1}{q_{n}}-\frac{1}{2}}\|W_{1}\|_{L^{n+1}}\|W_{2}\|_{L^{n+1}}.$
After summing (8) over a partition of unity, we obtain
(11) $(-\Delta_{g}-z)T(z)=\mathbf{1}+S(z),$
where $S(z)$ satisfies
(12) $\|S(z)W\|_{\mathfrak{S}^{2(n+1)}}\lesssim|z|^{\frac{1}{2q_{n}}}\|W\|_{L^{n+1}}.$
This follows from (9) and [8, Corollary 3] for $S_{1}(z)$, whereas for $S_{2}(z)$, (9) yields the stronger bound
$\|S_{2}(z)W\|_{\mathfrak{S}^{2(n+1)}}\leq\|S_{2}(z)W\|_{\mathfrak{S}^{2}}\lesssim|z|^{-\frac{1}{2}}\|W\|_{L^{2}}\lesssim|z|^{-\frac{1}{2}}\|W\|_{L^{n+1}}.$
In the last estimate, we used Hölder’s inequality and the fact that $M$ is compact.
Note that, even though $T(z)$ and $S(z)$ are not symmetric, their adjoints satisfy the same bounds (10), (12). Moreover, these estimates remain true when replacing $z$ by $\overline{z}$. Applying the resolvent operator $(-\Delta_{g}-z)^{-1}$ to (11) yields
(13) $W_{1}(-\Delta_{g}-z)^{-1}W_{2}=W_{1}T(z)W_{2}+W_{1}(-\Delta_{g}-z)^{-1}S(z)W_{2}$
and similarly for $z$ replaced by $\overline{z}$. We use the noncommutative Hölder inequality [16, Theorem 2.8] and (12) to bound
$\|W_{1}(-\Delta_{g}-z)^{-1}S(z)W_{2}\|_{\mathfrak{S}^{n+1}}$ $\leq\|W_{1}(-\Delta_{g}-z)^{-1}\|_{\mathfrak{S}^{2(n+1)}}\|S(z)W_{2}\|_{\mathfrak{S}^{2(n+1)}}$
(14) $\lesssim|z|^{\frac{1}{2q_{n}}}\|W_{1}(-\Delta_{g}-z)^{-1}\|_{\mathfrak{S}^{2(n+1)}}\|W_{2}\|_{L^{n+1}}.$
Then we use the resolvent identity
$(-\Delta_{g}-z)^{-1}(-\Delta_{g}-\overline{z})^{-1}=\frac{1}{2\mathrm{Im}\,z}((-\Delta_{g}-z)^{-1}-(-\Delta_{g}-\overline{z})^{-1})$
and the fact that $|\mathrm{Im}\,z|\gtrsim|z|^{\frac{1}{2}}$ in the considered range to obtain
$\|W_{1}(-\Delta_{g}-z)^{-1}\|_{\mathfrak{S}^{2(n+1)}}=\|W_{1}(-\Delta_{g}-z)^{-1}(-\Delta_{g}-\overline{z})^{-1}\overline{W_{1}}\|_{\mathfrak{S}^{n+1}}^{1/2}$
(15) $\lesssim|z|^{-\frac{1}{4}}(\|W_{1}(-\Delta_{g}-z)^{-1}\overline{W_{1}}\|_{\mathfrak{S}^{n+1}}+\|W_{1}(-\Delta_{g}-\overline{z})^{-1}\overline{W_{1}}\|_{\mathfrak{S}^{n+1}})^{1/2}.$
Combining (14)–(15) and using that $ab\lesssim\epsilon a^{2}+\epsilon^{-1}b^{2}$ for arbitrary $\epsilon>0$, we get
$\|W_{1}(-\Delta_{g}-z)^{-1}S(z)W_{2}\|_{\mathfrak{S}^{n+1}}$ $\lesssim\epsilon^{-1}|z|^{-\frac{1}{2}}(\|W_{1}(-\Delta_{g}-z)^{-1}\overline{W_{1}}\|_{\mathfrak{S}^{n+1}}$
(16) $+\|W_{1}(-\Delta_{g}-\overline{z})^{-1}\overline{W_{1}}\|_{\mathfrak{S}^{n+1}})+\epsilon|z|^{\frac{1}{q_{n}}}\|W_{2}\|^{2}_{L^{n+1}}$
and similarly for $z$ replaced by $\overline{z}$. We choose $\epsilon=C|z|^{-\frac{1}{2}}$ for some sufficiently large constant $C$. Then, using (10), (12), (13), (16) and their analogues for $\overline{z}$, and recalling that $W_{2}=\overline{W_{1}}$, we can absorb the $\epsilon^{-1}|z|^{-\frac{1}{2}}$ term to get (4) in this case.
Turning to the general case, in view of (10), it remains to prove
$\|W_{1}(-\Delta_{g}-z)^{-1}S(z)W_{2}\|_{\mathfrak{S}^{n+1}}\lesssim|z|^{\frac{1}{q_{n}}-\frac{1}{2}}\|W_{1}\|_{L^{n+1}}\|W_{2}\|_{L^{n+1}}.$
By the same arguments as before, this follows from (12) and the bound (4) with $W_{2}=\overline{W_{1}}$, which we have already proved. ∎
Acknowledgements: The author thanks Rupert Frank, Mikko Salo, Jeff Galkowksi and Xiaoyan Su for useful discussions and correspondence. Support by the Engineering & Physical Sciences Research Council [grant number EP/X011488/1] is acknowledged.